After a sailboat had been on the water for 4 h, a change in the wind direction reduced the average speed of the boat by 2 mph. The entire distance sailed was 161 mi. The total time spent sailing was 9 h. How far did the sailboat travel in the first 4 h
Answer by ankor@dixie-net.com(22740)

After a sailboat had been on the water for 4 h, a change in the wind direction reduced the average speed of the boat by 2 mph.
The entire distance sailed was 161 mi.
The total time spent sailing was 9 h.
How far did the sailboat travel in the first 4 h
:
let s = the speed of the boat the first 4 hrs
then
(s-2) = the speed of the last 5 hrs
:
Write a distance equation; dist = time * speed
4s + 5(s-2) = 161
4s + 5s - 10 = 161
9s = 161 + 10
s = 171/9
s = 19 mph for the first 4 hrs
then
4 * 19 = 76 mi traveled in the first 4 hrs
:
:
Check this, find the distance traveled in the last 5 hrs
5(17) = 85 mi
total dist
85 + 76 = 161